OpenStudy (anonymous):

a weight lifter lifts 150N of weight when doin curls. if he lifts the weight a distance of 40 cm in each curl (rep) how many reps must he make in 1.0 minute in order to reach a pwer of 25 watts? please explain why using power=work/time and work=force x distance

OpenStudy (lilg132):

power = work/time 150N / 60 seconds = 2.5 watts 25 / 2.5 = 10 reps I think

OpenStudy (anonymous):

haha thanks

OpenStudy (lilg132):

dont forget to click good answer :p

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

i don't think that's right, a newton is not a unit of work?

OpenStudy (anonymous):

WHOOPSS 250N not 150 srry!!

OpenStudy (anonymous):

do you know the formula to compute energy from work and distance?

OpenStudy (anonymous):

sorry dont :(

OpenStudy (anonymous):

which textbook are you using?

OpenStudy (lilg132):

newton is not a unit of work what is?

OpenStudy (anonymous):

a joule

OpenStudy (anonymous):

work (in joules) = force (in newtons) times displacement (in meters)

OpenStudy (anonymous):

im not using a text book my teacher just gave me a sheet with a bunch of problems

OpenStudy (anonymous):

ok, you must have some reference though? something you're learning from? what did the teacher teach you from?

OpenStudy (anonymous):

anyway, given the formula I just gave you can you compute the work done with each rep?

OpenStudy (anonymous):

she just taught us srry

OpenStudy (anonymous):

ok thaanks

OpenStudy (anonymous):

we haven't gotten through the problem yet

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the first step is to compute the work done with each rep

OpenStudy (anonymous):

it's just force (in newtons) times distance (in meters)

OpenStudy (lilg132):

150 * 0.4?

OpenStudy (anonymous):

wouldnt it be 40?

OpenStudy (anonymous):

yeah (except, it turns out the problem said 250 N, absolutezero mistyped it hte first time)

OpenStudy (anonymous):

yeaa sorry

OpenStudy (lilg132):

oops yes

OpenStudy (anonymous):

ok absolute0 so what number do you get for joules per rep?

OpenStudy (lilg132):

250 * 0.4 = 100 joules

OpenStudy (anonymous):

right, now with 100 joules you can do the computation you started to do earlier -- 100 joules * x reps divided by 60 seconds = 25 watts

OpenStudy (anonymous):

250 x 40 = 10000 is wat my calculator told me

OpenStudy (anonymous):

wait how is it 0.4?

OpenStudy (anonymous):

40 is centimeters, absolute0 - you have to make sure to do conversion to meters

OpenStudy (anonymous):

ooooooo thank you

OpenStudy (lilg132):

100 * 25 = 2500 2500 / 60 = 41 reps?

OpenStudy (lilg132):

41.666667 or 42?

OpenStudy (anonymous):

i get 15

OpenStudy (anonymous):

100 joules / 60 seconds = 1.66667 watts doing 1 rep per 60 seconds

OpenStudy (anonymous):

25/1.66667 = 15

OpenStudy (anonymous):

google calculator is really useful here because it does dimensional analysis for you

OpenStudy (anonymous):

actually this does the whole thing: http://tinyurl.com/7uhqkws

OpenStudy (anonymous):

thank you!! i apreciate it ;)

OpenStudy (anonymous):

no problem, and if i could ask, because i'm interested - are you in an online school, or a physical school?

OpenStudy (anonymous):

im in a physical school why?

OpenStudy (anonymous):

but they don't give you a textbook? how are you supposed to study at home?

OpenStudy (anonymous):

well the teacher just showed us this on the white board and besides this was is an assignment

OpenStudy (anonymous):

but where are you supposed to ge tthe formulas from and stuff?

OpenStudy (anonymous):

the teacher taught us this using a video then further explaining it

OpenStudy (anonymous):

but you don't have any written reference?

OpenStudy (anonymous):

not that i know of sorry

OpenStudy (anonymous):

interesting ...

OpenStudy (anonymous):

well last year and this year we didnt do much textbook work